find the equation of the line that is perpendicular to the line x-3y-7=0 and passes through the point (9,-3)in standard form and general form

Found 2 solutions by lwsshak3, stanbon:
Answer by lwsshak3(11628)

find the equation of the line that is perpendicular to the line x-3y-7=0 and passes through the point (9,-3)in standard form and general form
***
Standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
Slope of given line=-3
Slope of line perpendicular to given line=negative reciprocal of slope of given line=-(1/-3)=1/3
Equation of perpendicular line:
y=x/3+b
solve for b using coordinates of given points line passes thru.
-3=9/3+b
b=-6
Equation:
Standard form: y=x/3-6
or
General form: x-3y=18

Answer by stanbon(75887)

find the equation of the line that is perpendicular to the line x-3y-7=0 and passes through the point (9,-3)in standard form and general form
----
Given Line DATA:
x - 3y - 7 = 0
----
3y = x-7
y = (1/3)x -(7/3)
slope = 1/3
=======================
New Line DATA:
Form:: y = mx + b
slope = -3
passes thru 9,-3
--------------
-3 = -3*9 + b
b = 24
-----
New Equation:
y = -3x + 24
====================
Cheers,
Stan H.
====================